Solve the equation. If an answer is an excluded value state it. Clear the fractions from the equation first.
3x/x-4 - 6/x^2-7x+12 = 2x/x-3

Answer by jorel1380(3719)

3x/(x-4)-6/(x²-7x+12)=2x/(x-3)
3x/(x-4)-6/(x-4)(x-3)=2x/(x-3)
3x(x-3)-6=2x(x-4)
3x²-9x-6=2x²-8x
x²-x-6=0
(x+2)(x-3)=0
x=-2 or 3
Since 2x/(x-3) is undefined for x=3, we get only x=-2 as an answer. ☺☺☺☺